Month: December 2012

LA Area Legacy DSL DHCP Outage

A DHCP server has failed that serves a subset of legacy dsl customers in the LA area. We are currently working on re-routing its DHCP traffic to a backup server.

– Robbie

Update: Traffic has been re-routed and affected customers should be receiving new leases at this point.

Closure for Staff Party will be closed Friday from noon to five pm for our annual staff holiday party. Technical support will re-open 5pm-11pm, and we will be returning calls to anyone who has left a message during the closure. Our 24×7 network operations team will be available on call on the NOC hotline for our Enterprise dedicated circuit and colocation customers.

Database Cluster Maintenance

System Operations will be replacing a failed battery backup unit in one of our customer-facing database servers at 11:00pm PST on Thursday, December 20th, 2012. During this time, MySQL 5.x databases will be unavailable. The work is expected to take no more than 30 minutes.


** Update **

The maintenance window will me moved forward by 1 hour to 11:59pm tonight, December 20th.


** Update **

Database server is back up as of 12:30am.

— Joe (System Operations)

LATA9 ATM Maintenance

Tonight, December 19, starting at 11:59pm, we will be performing intrusive maintenance on legacy DSL and FRATM circuits in the Modesto/Stockton area. Expected downtime is 15 minutes.

-Tomoc and Robbie

We have encountered some issues during this maintenance window and still have some customers offline. We apologize for the inconvenience  and are working towards a solution asap. We will update this posting when the problem has been resolved.

Update: All remaining issues have been cleared up at this point. Please contact our support department as usual if you are having any further trouble with your DSL connection.

Update: A database error caused extended downtime for a subset of affected customers. All service should be restored at this point, so please contact support if you are still having trouble.

Upstream routing issue

At about 3:25PM today we detected a routing issue with an upstream provider that was causing network reachability issues for our users. We have rerouted around the issue and are closely monitoring the situation.  As of 3:35PM all connectivity should be restored.

-Sonic NOC

707.522.1000 returning all circuits busy

At 9am this morning 707.522.1000 began returning all circuits busy.  We are working with our vendor in order to remedy this issue.  In the mean time please use 1-888-766-4233 as an alternative.  All numbers to reach support directly remain unaffected.


–update:  707.522.1000 is currently up and working correctly. We will continue to monitor the situation.



– Jasper C. and Kavan Support

Support Contact Numbers Down

The customer contact lines 1-888-766-4233 and 707-522-1000 are down due to an error with our trunk provider. We are working diligently to bring both numbers back up. Customer Support can be reached at 707-547-3400.

— Jasper and Kim

Update: Service to 1-888-766-4233 has been restored.

Update: All lines are now restored.